Order: 50,000 units of penicillin G potassium/kg body mass/day, to be administered in for equally divided doses. penicillin G potassium 500,000 unit tabs are available. how many tabs will a 176 lb patient receive per dose?
Answer #1

Mass of the patient in kg = mass in lb / 2.20 = 176 / 2.2 = 80 kg

Units needed per patient = 50000 * mass of patient = 50000 * 80 = 4,000,000 units

Number of tabs = dosage per day / units per tab = 4,000,000 / 500000 = 8 tabs